a fluid is moving at 6.75mi/h in a section of a pipe of radius r = 20 cm. if the radius in another section is r = 0.35 m what is the flow speed there?
Answer #1

9. 5 – 6-ми) = 3 • 02 / s. я, – 20 см = p- 2 ) Це – ля, с - о.26 m _P, 2 аш а а , то SS-ѕесно и оил и о6 — - A2= area of cro